How much of a 40% antifreeze solution must a mechanic mix with an 80% antifreeze solution if 20 gallons of a 50% antifreeze solution are needed?
Here's what I tried:
Let x= amount of 40%, y= amount of 80%
0.40x + 0.80y = 0.5(20)
x + y = 20

Multiplied by 10 to get rid of decimals in first equation:
4x + 8y = 10
x + y = 20
Put into matrices and got:
(37.5, -17.5)
The answer is supposed to equal to 15 gal of 40%, 5 gal of 80% according to the lab worksheet I'm working on. I don't know what I am doing wrong.

Answer by ewatrrr(10682)

 
Hi
Let x= amount of 40%, y= amount of 80%
0.40x + 0.80y = 0.5(20)
x + y = 20 OR x = 20-y
good Work! substituting (20-y) for x
0.40(20-y) + 0.80y = 0.5(20)
.40y = .5(20)-.40(20)
y = .10(20)/.40
y = 5 gal of the 80% solution and 15gal of the 40% solution (5 + 15 = 20)